# Is Sarcasm Detection A Step-by-Step Reasoning Process in Large Language Models?

### v1

@@ Abstract
Elaborating a series of intermediate reasoning steps significantly improves the ability of large language models (LLMs) to solve complex problems, as such steps would evoke LLMs to think sequentially. However, human sarcasm understanding is often considered an intuitive and holistic cognitive process, in which various linguistic, contextual, and emotional cues are integrated to form a comprehensive understanding of the speaker’s true intention, which is argued not be limited to a step-by-step reasoning process. To verify this argument, we introduce a new prompting framework called SarcasmCue, which contains four prompting strategies, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC), which elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods. Through a comprehensive empirical comparison on four benchmarking datasets, we show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a considerable margin, and non-sequential prompting generally outperforms sequential prompting.
@@ 1 Introduction
Sarcasm is a subtle linguistic phenomenon that uses rhetorical devices such as hyperbole and figuration to convey true sentiments and intentions that are opposite to the literal meanings of the words used Wen et al. (2023); Zhang et al. (2023b). Sarcasm detection aims to combine different types of cues, such as linguistic features, contextual information, emotional knowledge, to form a comprehensive understanding of the author’s sarcastic attitude. Owing to its inherent ambivalence and figurative nature, sarcasm detection has persistently proven a formidable challenge spanning the eras from feature engineering to prompt engineering Yue et al. (2023); Zhang et al. (2023a).
Recent large language models have demonstrated impressive performance in downstream natural language processing (NLP) tasks, in which “System 1” - the fast, unconscious, and intuitive tasks, e.g., sentiment classification, topic analysis, etc., have been argued to be successfully performed Cui et al. (2024). Instead, increasing efforts have been devoted to the other class of tasks - “System 2”, which requires slow, deliberative and multi-step thinking, such as logical, mathematical, and commonsense reasoning tasks Wei et al. (2022). To improve the ability of LLMs to solve such complex problems, a widely adopted technique is to decompose complex problems into a series of intermediate solution steps prior to answer generation, and elicit LLMs to think step-by-step, such as chain of thought (CoT) Wei et al. (2022), tree of thought (ToT) Yao et al. (2024), graph of thought (GoT) Besta et al. (2024), etc.
However, sarcasm detection, as a holistic, intuitive, and non-rational cognitive process, is arguably in noncompliance with step-by-step logical reasoning due to two main reasons: (1) sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$). For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”; (2) sarcasm judgment is typically a fluid combination of various cues, where each cue holds equal importance to the judgment of sarcasm, and there is no rigid sequence of steps among them. As shown in Fig. 1, linguistic, contextual and emotional factors are all crucial for rendering the sentence as sarcastic. Hence, the main research question can be summarized as:
@@
To answer this question, we propose a theoretical framework, called SarcasmCue, based on the sequential and non-sequential prompting paradigm. It consists of four prompting methods, i.e., chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC). A cue is similar to a thought, which is concretely a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm, such as rhetorical devices, emotional words, etc. Each of the four prompting methods has its own focus and advantages. Specifically,
@@
  • CoC. It builds upon CoT prompting and harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention). It aims to: (1) identify the literal meaning and surface sentiment by extracting keywords, sentimental phrases, etc.; (2) deduce the true intention by scrutinizing special punctuation, rhetorical devices, cultural background, etc.; and (3) determine the inconsistency between surface sentiment and true intention. It has a typical linear structure.
@@
  • GoC. Generalizing over CoC, GoC frames the problem of sarcasm detection as a search over a graph and treats various cues (e.g., linguistic, contextual, emotional cues, etc.) as nodes, with the relations across cues represented as edges. Different from CoC and ToT, It allows language models to flexibly choose and weigh multiple cues when detecting sarcasm, rather than following a fixed hierarchy or linear reasoning path, unconstrained by the need for unique predecessor nodes. It represents a graphical structure. In summary, both CoC and GoC follow a step-by-step reasoning process.
@@
  • BoC. In contrast, BoC and ToC are proposed based on the assumption that sarcasm detection is not a step-by-step reasoning process. BoC is a bagging approach that constructs a pool of diverse cues and creates multiple cue subsets through randomly sampling q cues at each round. LLMs are employed to generate multiple predictions based on these subsets, and such predictions are aggregated to produce the final result via majority voting. It has a set-based structure.
@@
  • ToC. ToC treats each type of cues (namely linguistic, contextual, and emotional cues) as an independent, orthogonal view for sarcasm understanding and constructs a multi-view representation through the tensor product of these three types of cues. It allows language models to leverage higher-order interactions among the cues. ToC can be visualized as a 3D volumetric structure, where each coordinate axis corresponds to a distinct type of cue. This tensorial method aims to offer a more comprehensive and expressive means of fusing diverse cues.
We present empirical evaluations of the proposed prompting approaches across four sarcasm detection benchmarks over 2 SOTA LLMs (i.e., GPT-4o, LLaMA 3-8B), and compare their results against 3 SOTA prompting approaches (i.e., standard IO prompting, CoT, ToT). We show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a margin of 2%, and non-sequential prompting generally outperforms sequential prompting. Between the two LLMs, GPT-4o consistently beats LLaMA by a striking margin across all tasks.
The main contributions are concluded as follows:
@@
  • Our work is the first to investigate the step-wise nature of sarcasm judgment by using both sequential and non-sequential prompting methods.
  • We propose a new prompting framework that consists of four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC).
@@
  • Comprehensive experiments over four datasets demonstrate the superiority of the proposed prompting framework in zero-shot sarcasm detection.
@@ 2 Related Work
This section reviews two lines of research that form the basis of this work: CoT prompting and sarcasm detection.
@@ 2.1 Chain-of-Thought Prompting
Inspired by the step-by-step thinking ability of humans, CoT prompting was proposed to “prompt” language models to produce intermediate reasoning steps that lead to the final answer. Wei et al. Wei et al. (2022) made a formal definition of CoT prompting in LLMs and proved its effectiveness by presenting empirical evaluations on arithmetic reasoning benchmarks. This work pioneered the use of CoT prompting in NLP. However, its performance hinged on the quality of manually crafted prompts, which was a costly and unstable process. To fill this gap, Auto-CoT was proposed to automatically construct demonstrations with questions and reasoning chains Zhang et al. (2022). Different from Auto-CoT, Diao et al. Diao et al. (2023) presented an Active-Prompt approach to determine which questions were the most important and helpful to annotate from a pool of task-specific queries, for reducing the human engineering workload. The impressive results of CoT prompting have sparked a surge of exploration into designing CoT prompting strategies across various tasks Fei et al. (2023); Li et al. (2023); Zheng et al. (2023). For instance, Wang et al. Wang et al. (2024) used formal grammars as the intermediate reasoning steps for domain-specific language generation.
Furthermore, Yao et al. Yao et al. (2024) introduced a non-chain prompting framework, namely ToT, which made LLMs consider multiple different reasoning paths and self-evaluated choices to decide the next course of action. They proved the effectiveness of the ToT approach on the tasks requiring non-trivial planning or search. Beyond CoT and ToT approaches, Besta et al. Besta et al. (2024) modeled the information generated by an LLM as an arbitrary graph (i.e., GoT), where units of information were considered as vertices and the dependencies between these vertices were edges. Although the above-mentioned approaches have shown exceptional performance on various arithmetic and logical reasoning tasks, all of them adopt the sequential decoding paradigm of “let LLMs think step by step”. Contrarily, it is argued that sarcasm judgment does not conform to step-by-step logical reasoning, and there is a need to develop non-sequential prompting approaches.
@@ 2.2 Sarcasm Detection
Sarcasm detection is habitually treated as a text classification task, where the target is to identify whether the given text is sarcastic or not Zhang et al. (2024). It has evolved from early rule based and statistical learning based approaches to traditional neural methods, such as CNN, RNN, and further advanced to modern neural methods epitomized by Transformer models. In early stage, the rule based approaches infer the overall sarcasm polarity based on the refined sarcasm rules, such as the occurrence of the interjection word Zhang et al. (2023a). Statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, RF, NB, etc., to extract patterns and relationships within the data Zhou et al. (2023).
As deep learning based architectures have shown the superiority over statistical learning, numerous base neural networks, e.g., such as CNN Jain et al. (2020), LSTM Ghosh et al. (2018), GCN Liang et al. (2022), etc., have been predominantly utilized during the middle stage of sarcasm detection research, aiming to learn and extract complex features in an end-to-end fashion. As the field of deep learning continues to evolve, sarcasm detection research has stepped into the era of pre-trained language models (PLMs). An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations. For example, Liu et al. Liu et al. (2022) proposed a dual-channel framework by modeling both literal and implied sentiments separately. They also constructed two conflict prompts to elicit PLMs to generate the sarcasm polarity Liu et al. (2023b). Qiao et al. Qiao et al. (2023) presented a mutual-enhanced incongruity learning network to take advantage of the underlying consistency between the two modules to boost the performance. Tian et al. Tian et al. (2023) proposed a dynamic routing Transformer network to activate different routing transformer modules for modeling the dynamic mechanism in sarcasm detection.
However, the above-mentioned works still focus on how to utilize PLMs to extract effective features, without leveraging the extraodinary context learning capabilities of LLMs. In contrast, this paper makes the first attempt to explore the potential of prompting LLMs in sarcasm detection.
@@ 3 The Proposed Framework: SarcasmCue
The overall schematic illustration of the proposed SarcasmCue framework is illustrated in Fig. 2. We qualitatively compare SarcasmCue to other prompting approaches in Tab. 1. SarcasmCue is the only one to fully support chain-based, tree-based, graph-based, set-based and multidimensional array-based reasoning. It is also the only one that simultaneously supports both sequential and non-sequential prompting methods.
@@
Chain? Tree? Grap? Set? Tensor?
@@ 3.1 Task Definition
Consider a sarcasm detection task. Given the data set $\mathcal{D}=\left\{\left(\mathcal{X},\mathcal{Y}\right)\right\}$, where $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the input text sequence and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$ denotes the output label sequence. We use $\mathcal{L}_{\theta}$ to represent a large language model with parameter $\theta$. Our task is to leverage a collection of cues $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$ to bridge the input $\mathcal{X}$ and the output $\mathcal{Y}$, where each cue $c_{i}$ is a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm.
@@ 3.2 Chain of Contradiction
We capture the inherent paradoxical nature of sarcasm, which is the incongruity between the surface sentiment and the true intention, and introduce chain of contradiction, a CoT-style paradigm that allows LLMs to decompose the problem of sarcasm detection into intermediate steps and solve each before making decision (Fig. 2 (a)). Each cue $c_{k}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{k}|\mathcal{X},c_{1},c_{2},...,c_{k-1}\right)$ is sampled sequentially, then the output $\mathcal{Y}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|\mathcal{X},c_{1},...,c_{k}\right)$. A specific instantiation of CoC involves three steps:
@@
The output sequence $y_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1}\right)$ is generated from the language model $\mathcal{L}_{\theta}^{CoC}$ conditioned on input prompt $p_{1}$.
Step 2. We then ask LLM to carefully discover the true intention via the following prompt $p_{2}$:
Deduce what the sentence really means, namely the TRUE intention, by carefully checking any rhetorical devices, language style, unusual punctuation, common senses.
The output sequence, denoted as $y_{2}$, is generated from the language model conditioned on prompt $p_{2}$ as well as the previous interaction $p_{1},y_{1}$, formulated as $y_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2}\right).$
Step 3. We finally ask LLM to examine the consistency between surface sentiment and true intention and make the final prediction:
@@
$y_{3}$ is therefore generated based on a joint understanding of the preceding context $y_{1}$, $y_{2}$ and $p_{1}$, $p_{2}$, $p_{3}$: $y_{3}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2},y_{2},p_{3}\right)$. The sarcasm label is identified from $y_{3}$ as the output of CoC.
Notably, CoC is built based on the presumption that all the cues are linearly correlated, and detects human sarcasm through step-by-step reasoning. Different from the original CoT, however, the steps are explicitly designed for the sarcasm detection context. Further details are presented in Algorithm 1 in App. A.

### v2

@@ Abstract
Elaborating a series of intermediate reasoning steps significantly improves the ability of large language models (LLMs) to solve complex problems, as such steps would evoke LLMs to think sequentially. However, human sarcasm understanding is often considered an intuitive and holistic cognitive process, in which various linguistic, contextual, and emotional cues are integrated to form a comprehensive understanding, in a way that does not necessarily follow a step-by-step fashion. To verify the validity of this argument, we introduce a new prompting framework (called SarcasmCue) containing four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC), which elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods. Through a comprehensive empirical comparison on four benchmarks, we highlight three key findings: (1) CoC and GoC show superior performance with more advanced models like GPT-4 and Claude 3.5, with an improvement of 3.5% $\uparrow$. (2) ToC significantly outperforms other methods when smaller LLMs are evaluated, boosting the F1 score by 29.7% $\uparrow$ over the best baseline. (3) Our proposed framework consistently pushes the state-of-the-art (i.e., ToT) by 4.2%, 2.0%, 29.7%, and 58.2% in F1 scores across four datasets. This demonstrates the effectiveness and stability of the proposed framework. [^1] Our codes are available at https://github.com/qiuchili/llm_sarcasm_detection.
@@ 1 Introduction
Recent large language models have demonstrated impressive performance across downstream natural language processing (NLP) tasks, in which “System 1” - the fast, unconscious, and intuitive tasks, e.g., sentiment classification, topic analysis, etc., have been argued to be successfully performed Cui et al. (2024). Instead, increasing efforts have been devoted to the other class of tasks - “System 2”, which requires slow, deliberative and multi-steps thinking, such as logical, mathematical, and commonsense reasoning tasks Wei et al. (2022). To improve the ability of LLMs to solve such complex problems, a popular paradigm is to decompose complex problems into a series of intermediate solution steps, and elicit LLMs to think step-by-step, such as chain of thought (CoT) Wei et al. (2022), tree of thought (ToT) Yao et al. (2024), graph of thought (GoT) Besta et al. (2024), etc.
However, due to its inherent ambivalence and figurative nature, sarcasm detection is often considered a holistic and non-rational cognitive process that does not conform to step-by-step logical reasoning for two main reasons: (1) sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$). For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”; (2) sarcasm judgment is often considered a fluid combination of various cues. Each cue holds equal importance and there is no rigid sequence of steps among them, as shown in Fig. 1. Hence, the main research question can be summarized as:
@@
To answer this question, we propose a theoretical framework, called SarcasmCue, based on the sequential and non-sequential prompting paradigm. It consists of four prompting methods, i.e., chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC). Each method has its own focus and advantages. In this work, cue is similar to thought, being a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator for identifying sarcasm, such as rhetorical devices or emotional words. More specifically,
@@
  • CoC. It harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention). It aims to: (1) identify the surface sentiment by extracting keywords, etc.; (2) deduce the true intention by scrutinizing rhetorical devices, etc.; and (3) determine the inconsistency between them. It is a typical linear structure.
@@
  • GoC. Generalizing over CoC, GoC frames the problem of sarcasm detection as a search over a graph and treats various cues as nodes, with the relations across cues represented as edges. Unlike CoC and ToT, it goes beyond following a fixed hierarchy or linear reasoning path. In summary, both CoC and GoC follow the step-by-step reasoning process.
@@
  • BoC. BoC is a bagging approach that constructs a pool of diverse cues and randomly sampling multiple cue subsets. LLMs are employed to generate multiple predictions based on these subsets, and such predictions are aggregated to produce the final result. It is a set-based structure.
@@
  • ToC. ToC treats each type of cues (namely linguistic, contextual, and emotional cues) as an independent, orthogonal view for sarcasm understanding and constructs a multi-view representation through the tensor product. It allows language models to leverage higher-order interactions among the cues. ToC can be visualized as a 3D volumetric structure. Hence, BoC and ToC are proposed based on the assumption that sarcasm detection is not a step-by-step reasoning process.
  • Their correlation. These four methods represent an evolution from linear to nonlinear, and from a single perspective to multiple perspectives, together forming a comprehensive theoretical framework (SarcasmCue). Their design aims to adapt to various sarcasm detection scenarios.
We present empirical evaluations of the proposed prompting approaches across four benchmarks over 4 SOTA LLMs (i.e., GPT-4o, Claude 3.5 Sonnet, Llama 3-8B, Qwen 2-7B), and compare their results against 3 SOTA prompting approaches (i.e., standard IO prompting, CoT and ToT). we highlight three key observations: (1) When the base model is more advanced (such as GPT-4 and Claude 3.5 Sonnet), CoC and GoC show superior performance against the state-of-the-art (SoTA) baseline with an improvement of 3.5% $\uparrow$. (2) ToC achieves the best performance when smaller LLMs are evaluated. For example, in Llama 3-8B, ToC’s average F1 score of 65.24 represents a 29.7% improvement over the best baseline method, ToT. In Qwen 2-7B, ToC shows a 58.2% improvement over the best baseline method, IO. (3) Our proposed framework consistently pushes SoTA by 4.2%, 2.0%, 29.7% and 58.2% in F1 scores across four datasets. This demonstrates the effectiveness of the proposed framework. The main contributions are concluded as follows:
@@
  • Our work is the first to investigate the stepwise reasoning nature of sarcasm detection by using both sequential and non-sequential prompting methods.
  • We propose a new prompting framework that consists of four sub-methods, $viz.$ CoC, GoC, BoC and ToC.
@@
  • Comprehensive experiments over four datasets demonstrate the superiority of the proposed prompting framework.
@@ 2.1 Chain-of-Thought Prompting
Inspired by the step-by-step thinking ability of humans, CoT prompting was proposed to “prompt” language models to produce intermediate reasoning steps. Wei et al. Wei et al. (2022) made a formal definition of CoT prompting in LLMs and proved its effectiveness by presenting empirical evaluations on arithmetic reasoning benchmarks. However, its performance hinged on the quality of manually crafted prompts. To fill this gap, Auto-CoT was proposed to automatically construct demonstrations with questions and reasoning chains Zhang et al. (2022). Furthermore, Yao et al. Yao et al. (2024) introduced a non-chain prompting framework, namely ToT, which made LLMs consider multiple different reasoning paths to decide the next course of action. Beyond CoT and ToT approaches, Besta et al. Besta et al. (2024) modeled the information generated by an LLM as an arbitrary graph (i.e., GoT), where units of information were considered as vertices and the dependencies between these vertices were edges.
However, all of them adopt the sequential decoding paradigm of “let LLMs think step by step”. Contrarily, it is argued that sarcasm judgment does not conform to step-by-step logical reasoning, and there is an urgent need to develop non-sequential prompting approaches.
@@ 2.2 Sarcasm Detection
Sarcasm detection has evolved from early statistical learning based approaches to traditional neural methods, and further advanced to modern neural methods epitomized by Transformer models. In early stage, statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, NB, etc., to extract patterns and relationships within the data Zhang et al. (2023). As deep learning based architectures have shown the superiority, numerous base neural networks, e.g., such as CNN Jain et al. (2020), LSTM Ghosh et al. (2018), GCN Liang et al. (2022), etc., have been predominantly utilized during the middle stage of sarcasm detection research. Now, sarcasm detection research has stepped into the era of pre-trained language models (PLMs). An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations Liu et al. (2023).
Different from them, we propose four prompting methods to make the first attempt to explore the potential of prompting LLMs in sarcasm detection.
@@ 3 The Proposed Framework: SarcasmCue
The proposed SarcasmCue framework is illustrated in Fig. 2. We qualitatively compare SarcasmCue with other prompting approaches in Tab. 1. SarcasmCue is the only one to fully support chain-based, tree-based, graph-based, set-based and multidimensional array-based reasoning. It is also the only one that simultaneously supports both sequential and non-sequential prompting methods.
@@
Chain? Tree? Graph? Set? Tensor?
@@ 3.1 Task Definition
Given the data set $\mathcal{D}=\left\{\left(\mathcal{X},\mathcal{Y}\right)\right\}$, where $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the input text sequence and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$ denotes the output label sequence. We use $\mathcal{L}_{\theta}$ to represent a large language model with parameter $\theta$. Our task is to leverage a collection of cues $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$ to brige the input $\mathcal{X}$ and the output $\mathcal{Y}$, where each cue $c_{i}$ is a coherent language sequence that serves as an intermediate indicator toward identifying sarcasm.
@@ 3.2 Chain of Contradiction
We capture the inherent paradoxical nature of sarcasm, which is the incongruity between the surface sentiment and the true intention, and propose chain of contradiction, a CoT-style paradigm that allows LLMs to decompose the problem of sarcasm detection into intermediate steps and solve each before making decision (Fig. 2 (a)). Each cue $c_{k}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{k}|\mathcal{X},c_{1},c_{2},...,c_{k-1}\right)$ is sampled sequentially, then the output $\mathcal{Y}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|\mathcal{X},c_{1},...,c_{k}\right)$. A specific instantiation of CoC involves three steps:
@@
$c_{1}$ is the output sequence, which can be formulated as $c_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{1}|\mathcal{X},p_{1}\right)$.
Step 2. We thus ask LLM to carefully discover the true intention via the following prompt $p_{2}$:
Deduce what the sentence really means, namely the TRUE intention, by carefully checking any rhetorical devices, language style, unusual punctuations, common senses.
$c_{2}$ is the output sequence, which can be formulated as $c_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{2}|\mathcal{X},c_{1},p_{2}\right)$.
Step 3. Let LLM examine the consistency between surface sentiment and true intention and make the final prediction:
@@
CoC raises a presumption that the cues are linearly correlated, and detects human sarcasm through step-by-step reasoning. Further details see Algorithm 1 in App. A.